# StutterZero and StutterFormer: End-to-End Speech Conversion for Stuttering Transcription and Correction

### v1

@@ 2.5.2 StutterZero Spectrogram Decoder
The spectrogram decoder uses the previously predicted spectrogram and the current context vector as inputs. Two fully connected layers form a pre-net that compress and transform the previous spectrogram frames into a lower-dimensional representation. A well-trained pre-net simplifies the input and extracts the most salient features. If the raw spectrogram frames were used, the decoder might ”shortcut” the learning process by copying the previous frame or minimally modifying it [29] [30].
StutterZero employs a location-based attention mechanism, which is an extension of classical content-based attention. In content-based attention, the model computes an attention score between a query (representing the decoder’s current state) and a key (representing the encoder’s output at any time step). The score measures how relevant each encoder state is to the current decoding step. The attention weights are then obtained by normalizing these scores using a softmax function. These weights form a probability distribution that determines how much ”attention” the decoder should pay to each encoder state when predicting the next output. Finally, a weighted context vector is used for output prediction [31].
@@
Once the spectrogram decoder predicts a fluent spectrogram signal, the Griffin–Lim algorithm is used to reconstruct the phase data and generate an audio signal in the amplitude domain [32].
@@ 2.5.3 StutterZero Transcript Decoder
Previous work in text-to-speech and fluent speech conversion models has demonstrated the efficacy of multitask decoder architectures. Multitask training sums the loss for both the spectrogram and transcript decoders, creating a joint loss function that forces the training process to optimize the loss on both decoders [24] [20] [27].
@@ 2.5.4 StutterZero Training
The spectrogram decoder uses mean squared error (MSE) loss, where the model minimizes the difference between the ground-truth spectrogram and the predicted fluent spectrogram across all frequency bins and time steps [33, 34].
@@ 2.6 StutterFormer Model Architecture
StutterFormer maintains the same multitask architecture as StutterZero, but switches out internal layers for the Attention mechanism found in Transformers as described in [35]. Figure 4 displays a high-level summary of StutterFormer’s architecture.
@@ 2.6.1 Multi-Head Attention
Multi-head attention serves as the fundamental building blocks of StutterFormer. Compared to single-head attention, which computes attention in a single attention distribution, multi-head attention projects queries, keys, and values into multiple subspaces, applies attention in parallel, and then recombines the results. This makes it possible for the model to take a joint attention function to data from several learned subspaces. For instance, one head may focus on short-range syntactic dependencies only while another head tracks long-range semantic connections. This increase in flexibility allows multi-head attention to learn a greater breadth of information while keeping the parameter count relatively equal to a larger single-head attention module [35].
@@ 2.6.2 StutterFormer Encoder
The encoder input remains a log-Mel spectrogram. Because a Transformer processes all input frames in parallel, it does not have a sense of order by default. Sinusoidal positional encoding gives each time step a deterministic vector (based on the position index) for the embedding at each time step. Residual and layer normalization layers are used between the multi-head attention and feed-forward layers to mitigate vanishing or exploding gradients [36] [37]. Three multi-head attention units are utilized in the encoder, each consisting of four heads. Similarly to StutterZero, the encoder outputs a context vector – a learned hidden representation of the input.
@@ 2.6.3 StutterFormer Decoders
Finally, a post-net consisting of five 1-dimensional convolutions refines the predicted mel spectrogram to denoise and sharpen the final audio signal [29] [38].
@@ 2.6.4 StutterFormer Training
Like StutterZero, StutterFormer employs a hybrid loss function that combines multiple weighted losses. The spectrogram decoder also uses MSE loss between the predicted and ground truth spectrograms. In addition to MSE loss, the spectrogram decoder also computes a mean absolute error (MAE) loss. MAE loss measures the absolute difference of the predicted values and target values without squaring. When tested with spectrogram applications, MAE loss promotes sharper spectrograms that prevent the predicted fluent speech from sounding slurred [39]. The transcript decoder also uses cross-entropy loss.
StutterFormer is trained on a cosine annealing with warm restarts scheduler. This approach periodically ”restarts” the learning rate to help escape local minima and explore the loss landscape more effectively [40] [41] [42]. Like StutterZero, StutterFormer was trained for 1000 epochs.
@@ 3.1 Benchmarking
In addition to the ASR-based and end-to-end approaches used in this study, state-of-the-art fluent-speech ASR models were also evaluated to establish a baseline for how accurately current speech recognition systems transcribe stuttered speech. This study used the Whisper-Tiny, Whisper-Small, and Whisper-Medium pretrained fluent ASR models as baselines [24]. Any larger models such as Whisper-Large could not be tested due to memory and hardware limitations. The 10% validation data split was used to assess each of the six models.
Three metrics were used to evaluate all models: Word Error Rate (WER), Character Error Rate (CER), and BERTScore. CER is defined in Equation (3) as the proportion of incorrectly predicted characters compared to the ground truth string. To assess the semantic similarity between the ground truth and predicted utterances, a BERTScore is calculated using pre-trained contextual embeddings from the Bidirectional Encoder Representations from Transformers (BERT) language model [43]. Cosine similarity is used as a metric to quantify similarity between high-dimensional embedding vectors. The BERTScore defines how similar the two strings are semantically, meaning it is more forgiving towards minor transcription mistakes that still preserve the overall meaning of the utterance [43].
@@
ASR-based $04\pm 0.04 0.0$$02\pm 0.015$ $0.9516\pm 0.04$
StutterZero (end-to-end) $16\pm 0.013 0.110$$10\pm 0.096$ $0.9174\pm 0.034$
StutterFormer (end-to-end) $08\pm 0.07 0.0$$07\pm 0.011$ $0.9411\pm 0.012$
@@ 3.2 Statistical Significance
The non-parametric, two-sided Wilcoxon Signed-Rank Test is used to assess the statistical significance of the improvements achieved by this research compared to the baseline Whisper performance. This test is chosen due to its widespread use and established precedent in evaluating significant differences between models [44]. Comparing the performance of the ASR-based approach with Whisper-Medium (the best performing Whisper model), the Wilcoxson Test returns a test statistic of $77631.0$, a p-value of $1\mathrm{e}^{-100}$, significant at $\alpha=0.05$. Running the test comparing the performance of the end-to-end StutterZero and StutterFormer models against Whisper-Medium also returns a p-value of $1\mathrm{e}^{-100}$, significant at $\alpha=0.05$. This shows that both StutterZero and StutterFormer demonstrate a significant improvement over state-of-the-art fluent speech models.
@@ 3.3 Ablation Study
StutterZero (with transcript decoder) $16\pm 0.013 0.110$$10\pm 0.096$ $0.9174\pm 0.034$
Ablated StutterZero (without transcript decoder) $0.339\pm 0.106$ $0.260\pm 0.116$ $0.437\pm 0.103$
StutterFormer (with transcript decoder) $08\pm 0.07 0.0$$07\pm 0.011$ $0.9411\pm 0.012$
Ablated StutterFormer (without transcript decoder) $0.221\pm 0.163$ $0.194\pm 0.186$ $0.552\pm 0.146$
@@ 4 Discussion
StutterZero is not far behind, with an 88% transcription accuracy on the combined validation set and 84% on the FluencyBank subset. The slight decrease in accuracy may be attributed to the Transformer architecture’s efficiency in modeling speech and audio sequences. Indeed, correcting stuttered speech is a sequence-to-sequence task, similar to machine translation and other language tasks where Transformers have shown exceptional performance [31].
The observed performance improvements resonate with trends in related domains, such as dysarthric and accented speech recognition, where multitask and end-to-end approaches have consistently demonstrated robustness to atypical input [21, 12]. By showing that stutter correction benefits strongly from the integration of textual and acoustic objectives, our findings suggest that linguistic supervision is not merely auxiliary but foundational. This echoes psycholinguistic accounts that disfluency cannot be treated as random noise but reflects systematic deviations in speech planning and motor execution. In this sense, our work challenges the assumption – common in early DSP and ASR pipelines – that acoustic correction can be divorced from lexical anchoring.
The superior performance of StutterFormer over StutterZero cannot be attributed solely to increased model capacity. Rather, Transformer attention provides specific advantages for disfluency correction. Multi-head self-attention captures long-range dependencies, allowing the model to flexibly relate repeated or prolonged segments of speech to their fluent counterparts. This mechanism facilitates alignment across syllables and words, which is essential when disfluencies span multiple phonemic units. In addition, attention layers can simultaneously model both global and local phonetic details, helping to preserve intonation and rhythm while removing interruptions. These properties explain why attention-based architectures are particularly well-suited for mapping stuttered to fluent speech [45].
@@ 4.1.2 Need for Increased Dataset Diversity
The diversity of the training corpus was constrained by limited access to datasets such as UCLASS or FluencyBank. The absence of these datasets restricted coverage of various speaker demographics, accents, and tones, which in turn limits generalizability. There may be risks of overfitting and undergeneralization when training on only the SEP-28K data from the LibriStutter datasets. Even though five-fold cross-validation and testing on a completely new dataset was used to produce a candid estimate of the true performance both approaches, future steps should focus on data augmentation and training on larger datasets. Additional experiments using diverse corpora such as UCLASS or AS-70 [46] are needed.
@@ 4.2.1 Prosody-aware Fine-tuning
To alleviate the prosodic loss due to training on TTS-generated fluent speech samples, using multiple encoders to capture the tonal and prosodic content of the stuttered speech could be explored. Multi-encoder models have been explored in dysarthic speech conversion, specifically using a prosody encoder to extract prosodic features [21]. This would enable fluent outputs that retain the speaker’s identity, pitch, and expressive nuance. Additionally, expanding to multilingual and low-resource languages through cross-lingual pretraining and transfer learning would extend accessibility to a wider global population [46].
@@ 4.2.2 Dataset Expansion and Multilinguality
The current study is limited by access to a small number of corpora. Expanding to diverse datasets such as UCLASS, FluencyBank, and AS-70 [46], as well as developing multilingual training resources, would enhance model generalizability. Cross-lingual pretraining and transfer learning offer promising strategies to extend accessibility to low-resource and global language communities, ensuring that stutter correction technology benefits a wider population.
@@ 4.2.3 Integration with Clinical Practices
StutterZero and StutterFormer have great potential in automating and assisting with delayed auditory feedback (DAF), a common technique used in speech-language therapy. It involves recording and playing back a PWS’s speech after a brief delay (usually a few milliseconds to fractions of a second) [47] [48]. Playback of the fluent speech can demonstrate how fluent speech is supposed to sound, thereby reinforcing self-monitoring and reducing disfluencies. Instead of the individual hearing their disfluent utterances delayed, the model could provide them with a fluent version of what they intended to say. This would supply the brain with consistent, fluent auditory feedback, potentially reducing the reinforcement of stuttered patterns while strengthening neural pathways associated with fluent production. Indeed, speaking in unison with a fluent signal (an external “fluent version” of one’s speech content) reliably induces near-instant fluency in most PWS [49].

### v2

@@ 2.5.2 StutterZero Spectrogram Decoder
The spectrogram decoder uses the previously predicted spectrogram and the current context vector as inputs. Two fully connected layers form a pre-net that compress and transform the previous spectrogram frames into a lower-dimensional representation. A well-trained pre-net simplifies the input and extracts the most salient features. If the raw spectrogram frames were used, the decoder might ”shortcut” the learning process by copying the previous frame or minimally modifying it [shen2018] [wang2017].
StutterZero employs a location-based attention mechanism, which is an extension of classical content-based attention. In content-based attention, the model computes an attention score between a query (representing the decoder’s current state) and a key (representing the encoder’s output at any time step). The score measures how relevant each encoder state is to the current decoding step. The attention weights are then obtained by normalizing these scores using a softmax function. These weights form a probability distribution that determines how much ”attention” the decoder should pay to each encoder state when predicting the next output. Finally, a weighted context vector is used for output prediction [chorowski2015].
@@
Once the spectrogram decoder predicts a fluent spectrogram signal, the Griffin–Lim algorithm is used to reconstruct the phase data and generate an audio signal in the amplitude domain [griffin1984].
@@ 2.5.3 StutterZero Transcript Decoder
Previous work in text-to-speech and fluent speech conversion models has demonstrated the efficacy of multitask decoder architectures. Multitask training sums the loss for both the spectrogram and transcript decoders, creating a joint loss function that forces the training process to optimize the loss on both decoders [radford2022] [toshniwal2017] [zhang2021].
@@ 2.5.4 StutterZero Training
The spectrogram decoder uses mean squared error (MSE) loss, where the model minimizes the difference between the ground-truth spectrogram and the predicted fluent spectrogram across all frequency bins and time steps [rafaely2025, wang2009].
@@ 2.6 StutterFormer Model Architecture
StutterFormer maintains the same multitask architecture as StutterZero, but switches out internal layers for the Attention mechanism found in Transformers as described in [vaswani2023]. Figure 4 displays a high-level summary of StutterFormer’s architecture.
@@ 2.6.1 Multi-Head Attention
Multi-head attention serves as the fundamental building blocks of StutterFormer. Compared to single-head attention, which computes attention in a single attention distribution, multi-head attention projects queries, keys, and values into multiple subspaces, applies attention in parallel, and then recombines the results. This makes it possible for the model to take a joint attention function to data from several learned subspaces. For instance, one head may focus on short-range syntactic dependencies only while another head tracks long-range semantic connections. This increase in flexibility allows multi-head attention to learn a greater breadth of information while keeping the parameter count relatively equal to a larger single-head attention module [vaswani2023].
@@ 2.6.2 StutterFormer Encoder
The encoder input remains a log-Mel spectrogram. Because a Transformer processes all input frames in parallel, it does not have a sense of order by default. Sinusoidal positional encoding gives each time step a deterministic vector (based on the position index) for the embedding at each time step. Residual and layer normalization layers are used between the multi-head attention and feed-forward layers to mitigate vanishing or exploding gradients [borawar2023] [he2015]. Three multi-head attention units are utilized in the encoder, each consisting of four heads. Similarly to StutterZero, the encoder outputs a context vector – a learned hidden representation of the input.
@@ 2.6.3 StutterFormer Decoders
Finally, a post-net consisting of five 1-dimensional convolutions refines the predicted mel spectrogram to denoise and sharpen the final audio signal [shen2018] [ren2019].
@@ 2.6.4 StutterFormer Training
Like StutterZero, StutterFormer employs a hybrid loss function that combines multiple weighted losses. The spectrogram decoder also uses MSE loss between the predicted and ground truth spectrograms. In addition to MSE loss, the spectrogram decoder also computes a mean absolute error (MAE) loss. MAE loss measures the absolute difference of the predicted values and target values without squaring. When tested with spectrogram applications, MAE loss promotes sharper spectrograms that prevent the predicted fluent speech from sounding slurred [guso2022]. The transcript decoder also uses cross-entropy loss.
StutterFormer is trained on a cosine annealing with warm restarts scheduler. This approach periodically ”restarts” the learning rate to help escape local minima and explore the loss landscape more effectively [liu2022] [jiang2024] [cazenave2022a]. Like StutterZero, StutterFormer was trained for 1000 epochs.
@@ 3.1 Benchmarking
In addition to the ASR-based and end-to-end approaches used in this study, state-of-the-art fluent-speech ASR models were also evaluated to establish a baseline for how accurately current speech recognition systems transcribe stuttered speech. This study used the Whisper-Tiny, Whisper-Small, and Whisper-Medium pretrained fluent ASR models as baselines [radford2022]. Any larger models such as Whisper-Large could not be tested due to memory and hardware limitations. The 10% validation data split was used to assess each of the six models.
Three metrics were used to evaluate all models: Word Error Rate (WER), Character Error Rate (CER), and BERTScore. CER is defined in Equation (3) as the proportion of incorrectly predicted characters compared to the ground truth string. To assess the semantic similarity between the ground truth and predicted utterances, a BERTScore is calculated using pre-trained contextual embeddings from the Bidirectional Encoder Representations from Transformers (BERT) language model [zhang2020]. Cosine similarity is used as a metric to quantify similarity between high-dimensional embedding vectors. The BERTScore defines how similar the two strings are semantically, meaning it is more forgiving towards minor transcription mistakes that still preserve the overall meaning of the utterance [zhang2020].
@@
ASR-based $04\pm 0.01 0.0$$02\pm 0.01$ $0.9516\pm 0.04$
StutterZero (end-to-end) $16\pm 0.013 0.110$$10\pm 0.052$ $0.9174\pm 0.034$
StutterFormer (end-to-end) $08\pm 0.03 0.0$$07\pm 0.011$ $0.9411\pm 0.012$
@@ 3.2 Statistical Significance
The non-parametric, two-sided Wilcoxon Signed-Rank Test is used to assess the statistical significance of the improvements achieved by this research compared to the baseline Whisper performance. This test is chosen due to its widespread use and established precedent in evaluating significant differences between models [rainio2024]. Comparing the performance of the ASR-based approach with Whisper-Medium (the best performing Whisper model), the Wilcoxson Test returns a test statistic of $77631.0$, a p-value of $1\mathrm{e}^{-100}$, significant at $\alpha=0.05$. Running the test comparing the performance of the end-to-end StutterZero and StutterFormer models against Whisper-Medium also returns a p-value of $1\mathrm{e}^{-100}$, significant at $\alpha=0.05$. This shows that both StutterZero and StutterFormer demonstrate a significant improvement over state-of-the-art fluent speech models.
@@ 3.3 Ablation Study
StutterZero (with transcript decoder) $16\pm 0.013 0.110$$10\pm 0.052$ $0.9174\pm 0.034$
Ablated StutterZero (without transcript decoder) $0.339\pm 0.011$ $0.260\pm 0.012$ $0.437\pm 0.010$
StutterFormer (with transcript decoder) $08\pm 0.03 0.0$$07\pm 0.011$ $0.9411\pm 0.012$
Ablated StutterFormer (without transcript decoder) $0.221\pm 0.016$ $0.194\pm 0.019$ $0.552\pm 0.015$
@@ 4 Discussion
StutterZero is not far behind, with an 88% transcription accuracy on the combined validation set and 84% on the FluencyBank subset. The slight decrease in accuracy may be attributed to the Transformer architecture’s efficiency in modeling speech and audio sequences. Indeed, correcting stuttered speech is a sequence-to-sequence task, similar to machine translation and other language tasks where Transformers have shown exceptional performance [chorowski2015].
The observed performance improvements resonate with trends in related domains, such as dysarthric and accented speech recognition, where multitask and end-to-end approaches have consistently demonstrated robustness to atypical input [wang2022, mujtaba2025]. By showing that stutter correction benefits strongly from the integration of textual and acoustic objectives, our findings suggest that linguistic supervision is not merely auxiliary but foundational. This echoes psycholinguistic accounts that disfluency cannot be treated as random noise but reflects systematic deviations in speech planning and motor execution. In this sense, our work challenges the assumption – common in early DSP and ASR pipelines – that acoustic correction can be divorced from lexical anchoring.
The superior performance of StutterFormer over StutterZero cannot be attributed solely to increased model capacity. Rather, Transformer attention provides specific advantages for disfluency correction. Multi-head self-attention captures long-range dependencies, allowing the model to flexibly relate repeated or prolonged segments of speech to their fluent counterparts. This mechanism facilitates alignment across syllables and words, which is essential when disfluencies span multiple phonemic units. In addition, attention layers can simultaneously model both global and local phonetic details, helping to preserve intonation and rhythm while removing interruptions. These properties explain why attention-based architectures are particularly well-suited for mapping stuttered to fluent speech [zeyer2019].
@@ 4.1.2 Need for Increased Dataset Diversity
The diversity of the training corpus was constrained by limited access to datasets such as UCLASS or FluencyBank. The absence of these datasets restricted coverage of various speaker demographics, accents, and tones, which in turn limits generalizability. There may be risks of overfitting and undergeneralization when training on only the SEP-28K data from the LibriStutter datasets. Even though five-fold cross-validation and testing on a completely new dataset was used to produce a candid estimate of the true performance both approaches, future steps should focus on data augmentation and training on larger datasets. Additional experiments using diverse corpora such as UCLASS or AS-70 [gong2024] are needed.
@@ 4.2.1 Prosody-aware Fine-tuning
To alleviate the prosodic loss due to training on TTS-generated fluent speech samples, using multiple encoders to capture the tonal and prosodic content of the stuttered speech could be explored. Multi-encoder models have been explored in dysarthic speech conversion, specifically using a prosody encoder to extract prosodic features [wang2022]. This would enable fluent outputs that retain the speaker’s identity, pitch, and expressive nuance. Additionally, expanding to multilingual and low-resource languages through cross-lingual pretraining and transfer learning would extend accessibility to a wider global population [gong2024].
@@ 4.2.2 Dataset Expansion and Multilinguality
The current study is limited by access to a small number of corpora. Expanding to diverse datasets such as UCLASS, FluencyBank, and AS-70 [gong2024], as well as developing multilingual training resources, would enhance model generalizability. Cross-lingual pretraining and transfer learning offer promising strategies to extend accessibility to low-resource and global language communities, ensuring that stutter correction technology benefits a wider population.
@@ 4.2.3 Integration with Clinical Practices
StutterZero and StutterFormer have great potential in automating and assisting with delayed auditory feedback (DAF), a common technique used in speech-language therapy. It involves recording and playing back a PWS’s speech after a brief delay (usually a few milliseconds to fractions of a second) [ozker2025] [buzzeti2018]. Playback of the fluent speech can demonstrate how fluent speech is supposed to sound, thereby reinforcing self-monitoring and reducing disfluencies. Instead of the individual hearing their disfluent utterances delayed, the model could provide them with a fluent version of what they intended to say. This would supply the brain with consistent, fluent auditory feedback, potentially reducing the reinforcement of stuttered patterns while strengthening neural pathways associated with fluent production. Indeed, speaking in unison with a fluent signal (an external “fluent version” of one’s speech content) reliably induces near-instant fluency in most PWS [kalinowski2003].